OpenStudy (anonymous):

A Ferris wheel has a radius of 20 m and turns once every 40 seconds. Passengers get on at the bottom, which is 1 m above the ground. what is the cos equation

OpenStudy (anonymous):

radius of 20means your amplitude will be 20

OpenStudy (anonymous):

turns once every 40 seconds means period is 40, so put \(40=\frac{2\pi}{b}\) solve for \(b\) get \(b=\frac{\pi}{20}\)

OpenStudy (anonymous):

I did the equation, i'm just not sure if its right. I got y= -2-cos(piT/20)+21\[y=\]

OpenStudy (anonymous):

-20 for a

OpenStudy (anonymous):

that looks exactly right

OpenStudy (anonymous):

\[-20\cos(\frac{\pi t}{20})+21\] i think does it

OpenStudy (anonymous):

really? thank you. do you mind if i ask you another question?

OpenStudy (anonymous):

sure i will see if i can help

OpenStudy (anonymous):

there is this question in my book and it gives a table with all the months 1-12, and then each month has the duration of daylight in (h:m). So month 1, has duration of 9:05, month 2, has duration of 9:57, month 3 has duration of 11:12. It asks to express the duration as decimal values to the nearest hundredth of an hour.

OpenStudy (anonymous):

how would i go that

OpenStudy (anonymous):

divide by 60

OpenStudy (anonymous):

\(\frac{.05}{60}\) for the first one, use a calculator

OpenStudy (anonymous):

the answer at the back says that for month 1, the decimal duration is 9.08

OpenStudy (anonymous):

for the 2ed its 9.95

OpenStudy (anonymous):

and for the 3rd its 11.2

OpenStudy (anonymous):

then of course add 9 so the first one is something like 9.0008

OpenStudy (anonymous):

the first one is 9.08

OpenStudy (anonymous):

ooops i must be tired

OpenStudy (anonymous):

divide by 60, multiply by 100

OpenStudy (anonymous):

i.e. divide by 60, move the decimal place over two

OpenStudy (anonymous):

i get 9.83

OpenStudy (anonymous):

ok let me stop being stupid divide the minutes by 60

OpenStudy (anonymous):

wait, its working

OpenStudy (anonymous):

so you have 9 hours and 5 minutes 5 minutes convert to decimal by \(\frac{5}{60}=.083\) rounded

OpenStudy (anonymous):

so first one is \(9.083\)

OpenStudy (anonymous):

i divide the min by 60 and multiply by 100 , yes

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

well what i said was silly just divide the minutes by 60 so for example \(\frac{57}{60}=.95\) and second one should be \(9.95\) etc yw

OpenStudy (anonymous):

its better than nothing. thank you so so much. could you explain by i'm dividing by 60 though?